The angle x through which a bicycle wheel turns is given by x = a + bt2 + ct3 where a, b, and c are constants.


Calculate the velocity and acceleration of the wheel as a function of time.


1
Expert's answer
2021-04-26T16:48:52-0400

"x = a+bt^2+ct^3"

"\text{The angular velocity of rotation is a vector numerically equal to the first derivative}"

"\text{of the angle of rotation of the body}"

"\vec\omega= x'=2bt+3ct^2"

"\text{Angular acceleration is a vector quantity equal to the first derivative}"

"\text{{of the angular velocity with respect to time}}"

"\vec\alpha=\vec\omega'=2b+6ct"

"\text{Answer : }\vec\omega= 2bt+3ct^2;\vec\alpha=2b+6ct"